Grade 12Mechanics

A bucket contains water filled upto a height 15 cm the bucket is tied to a rope which is passed over a frictionless light pulley and the other end of the rope is tied to a weight of mass which is half of that of the bucket+water. The water pressure above atmosphere pressure at the bottom is

Ritik
9 Years agoGrade 12

3 Answers

Jatin
9 Years ago
As the bucket will not move because mass hanging is half of that of water and bucket so there is no acceleration so pressure at bottom is As height is in centimeter15 centimeter=3/20 meterP(bottom)=P°(atmospheric)+ (3/20)*density*gSo pressure above atmosphere pressure is(3/20)*density*gDensity of water=1000Kg/litre or 1gm/mlAnswer=(3/20)*1000*10=1500 in pascal(Nm-2)If your answer have different units feel free to ask again and let me know if you have any other doubt
Ayush
7 Years ago
In the above ans my friend made a small mistake by not considering downward acceleration of (bucket+water). Hence getting a wrong solution.
Let m be the mass of (bucket + water)
According to question mass of body on other side of pulley will be ½m
Net force on system will be:
Fnet= mg-½mg = ½mg
Acceleration of the system will be :
Anet= Fnet/total mass of system
Or, Anet = ½mg/(m+½m) = g/3
Pressure at the bottom above the atmospheric pressure will be given by rho*g*h
Rho= density of water = 1000kg/m^³
g = 10m/s^² and h= 15 cm or, 3/20 m
 
Therefore , the required pressure is rho * g* h = 1000*10*3/20 = 1000 Pa or 1kPa.
 
Thank you.
Ayush
7 Years ago
Sorry I myself made a mistake.
Net acceleration downwards will be (g-a) = g-g/3 = 2g/3.
Pressure= 1000*(2/3)*10*(3/20) = 1000 Pa or 1kPa